OpenStudy (anonymous):

a boat travels 2mi south and then 8 mi east.how far is the boat from its starting point? a=3.5mi b=34 mi c=8.2 mi d=3.2mi

OpenStudy (anonymous):



OpenStudy (anonymous):

what law do you know of , when it comes to a right angled triangle?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

exactly!!

OpenStudy (anonymous):

so 2*2=4 8*8=64 64+4=68

OpenStudy (anonymous):

so what what would be the answers in those options

OpenStudy (anonymous):

c. square root of 68 is 8.24